# Logical perspectives on learning statistical objects

### v1

@@ 4 Hypothesis classes from statistical objects
Randomization and random variable-based classes. We can embed the distribution class and the dual distribution class in a more general family of hypothesis classes, where parameters and range elements are treated symmetrically. We present the construction for logic-based classes, as in (Ben Yaacov, 2009; 2013). But the construction can be adapted for any hypothesis class.
Definition 7 (Well behaved collection of $M$-valued random variables).
@@
Given a CL structure $M$, an $M$-valued random variables os functions from some atomless probability space $\Omega^{*}=(\Omega,\Sigma,\mu_{0})$ to $M$. A collection of $M$-valued random variables is well-behaved if for each CL formula $\phi(\vec{x})$ over $\mathfrak{C}$, and each tuple $\vec{X}$ of functions in the set, $\phi(\vec{X})$ is $\Omega^{*}$-measurable.
For example, the set of functions with countable image, such that the pre-image of each point is in $\Sigma$, is well-behaved. Thus, given $F_{1}\ldots F_{k}$ from well-behaved $S$, the composition of $\phi(F_{1}\ldots F_{k})$ is a real-valued random variable on $\Omega$.
Definition 8 (Randomization of a structure).
Consider a classical first-order structure or a continuous-valued structure $\mathfrak{C}$ with domain $M$. We can form a new structure, the randomization of $\mathfrak{C}$, ${\mathcal{R}AND}_{\mathfrak{C}}$ (Keisler, 1999; Ben Yaacov & Keisler, 2009; Ben Yaacov, 2013). For each (classical first-order or real-valued) formula $\phi(x_{1}\ldots x_{k})$ over the signature of $\mathfrak{C}$, there is a new real-valued predicate $E\phi(X_{1}\ldots X_{k})$, where $X_{1}\ldots X_{k}$ are new variables.
This completes the description of the signature. We now describe the structure. Its domain is a well-behaved subset $S$ of the $M$-valued random variables, based on atomless probability space $\Omega^{*}=(\Omega,\Sigma,\mu_{0})$. Given $F_{1}\ldots F_{k}$ from $S$, the composition of $\phi(F_{1}\ldots F_{k})$ is a real-valued random variable on $\Omega$, and we set $E\phi$ evaluated at $F_{1}\ldots F_{k}$ to be the expected value of this random variable under $\mu_{0}$.
The randomization is not unique. However, all the results presented will hold for any randomization, and thus we will sometimes abuse notation by referring to it as if it were unique.
Note that we consider the randomization of $\mathfrak{C}$ as a CL structure, and thus we can apply CL connectives and quantification to form new formulas based on the primitive formulas $E\phi$ above. Our requirement on well-behavedness says that our random variables “play well” with the expectation of formulas from the base structure. The following result states that if we have this, then these random variables also play well with CL formulas over the randomization:
Fact 9.
Ben Yaacov & Keisler (2009); Ben Yaacov (2013). For any real-valued formula $f(x)$ in the language of ${\mathcal{R}AND}_{\mathfrak{C}}$, any probability space $(\Omega,\mathcal{B},\mu)$, and any $a\in({\mathcal{R}AND}_{\mathfrak{C}})^{x}$, the function $\omega\mapsto f(a(\omega))$ is measurable.
Now fix a variable-partitioned formula $\phi(\vec{x};\vec{y})$ in the language of a structure $\mathfrak{M}$, and consider the real-valued formula $E\phi$ in the randomization. It has free variables $\vec{X}$ and $\vec{Y}$, both ranging over random variables. In (Ben Yaacov, 2009) (for concept classes) and (Ben Yaacov, 2013) (for real-valued based classes) it is shown that moving from $\phi$ in $\mathfrak{C}$ to $E\phi$ in the randomization preserves agnostic PAC learnability:
Fact 10.
(Ben Yaacov, 2013, Theorem 4.10) If the hypothesis class induced by the formula $\phi(\vec{x};\vec{y})$ over a continuous logic structure $\mathfrak{C}$ is agnostic PAC learnable, then so is the class corresponding to the real-valued formula $E\phi(\vec{X};\vec{Y})$ in the randomization. Further, if the structure $\mathfrak{C}$ is NIP – that is, every definable family is PAC learnable – then the same holds for the real-valued structure ${\mathcal{R}AND}_{\mathfrak{C}}$.
Fact 10 will also be a corollary of our Theorem 5, with the latter providing more precise bounds.
We now want to apply this to the distribution and dual distribution class. We need to address the fact that in these classes, we consider distributions on the parameter space or the range space, not on some external sample space. Given two measure spaces $(\Omega,\Sigma,\mu_{0})$ and $(X,\Sigma^{\prime},\mu^{\prime})$, and a function $F$ from $\Omega$ to $X$, we say that $F$ induces $\mu_{0}$ from $\mu^{\prime}$ if for every $S^{\prime}\in\Sigma^{\prime}$, $F^{-1}(S^{\prime})\in\Sigma$ and $\mu_{0}(F^{-1}(S^{\prime}))=\mu^{\prime}(S^{\prime})$. The following is a basic result in measure theory (Fremlin, 2002), see Appendix D:
Proposition 3.
For any set $X$, we can choose $(\Omega,\Sigma,\mu)$ and well-behaved set of random variables $RV_{X}$ such that for each probability measure $\mu^{\prime},\Sigma^{\prime}$ on $X$ there is $F\in RV_{X}$ inducing $\mu^{\prime}$ from $\mu$.
From now on for a model $\mathfrak{C}$ we assume a suitably rich probability space and a well-behaved set of random variables, and refer to the randomization structure based on them as ${\mathcal{R}AN}_{\mathfrak{C}}$. We use analogous terminology for a partitioned formula $\phi(\vec{x};\vec{y})$ over $\mathfrak{C}$ – by the randomization class of $\phi$ we refer to the hypothesis class corresponding to the CL formula $E\phi$. We could similarly refer to the randomization class for a hypothesis class ${\mathcal{H}}$ over range space $X$ and parameter space $P$, referring to the hypothesis class with random variables into $X$ as range elements and random variables into $P$ as parameters.
We start with the observation that agnostic learnability of the randomization classes is sufficient to get agnostic learnability for the distribution and dual distribution classes:
Proposition 4.
For a formula $\phi(x_{1}\ldots x_{j};y_{1}\ldots y_{k})$ over a model $\mathfrak{C}$, if the randomization class corresponding to the formula $E\phi(\vec{X};\vec{Y})$ is agnostic PAC learnable, then so are the distribution class and dual distribution class for ${\mathcal{H}}_{\phi}$. Similarly for a hypothesis class ${\mathcal{H}}$, if its randomization class is agnostic PAC learnable, then so are its distribution and dual distribution classes. The same holds for agnostic online learnability.
Proof.
We consider only the PAC case for simplicity.
Consider the functions in the randomization class $E\phi(\vec{X};\vec{Y})$, varying the random parameters $\vec{Y}$ but restricting the function arguments $\vec{X}$ to be deterministic. These are functions that map a vector $\vec{x}$ to the expectation of $\phi(\vec{x},\vec{Y})$ with respect to a given random variable $\vec{Y}$. By Proposition 3, the distributions induced by $\vec{Y}$ cover all distributions over the parameter space $M^{k}$. Thus this class of restrictions give exactly the distribution class over ${\mathcal{H}}_{\phi(\vec{x};\vec{y})}$. But if a class of functions is agnostic PAC learnable, then so is the class obtained by restricting its domain to a subset.
For the dual distribution class, we can apply Fact 7 and then the argument above. ∎
Example 4.2. Consider the base model to be additive arithmetic $(\mathbb{N},+,<)$. We can write a partitioned formula $\phi(x;y_{1},y_{2})$ that states that $x$ is an even integer between $y_{1}$ and $y_{2}$. Thus as we vary the parameters $y_{1},y_{2}$ we get the set of intervals intersected with the even numbers.
Fix a probability space $\Omega^{*}=(\Omega,\Sigma,\mu_{0})$ and consider the hypothesis space where the parameter space $Y$ consists of “pairs of random integers”: pairs of $\Omega^{*}$ measurable functions from $\Omega$ into $\mathbb{N}$ with countable range. Given a parameter $Y=\langle Y_{1},Y_{2}\rangle$, we can consider the function $h_{Y_{1},Y_{2}}$ mapping random integers $X$ ($\Omega^{*}$-measurable functions to $\mathbb{N}$) to the probability that the integer produced by $X$ is even and between $Y_{1},Y_{2}$: that is, we get the function class for $E\phi(X;Y_{1},Y_{2})$. If we restrict each of the functions in the class to the deterministic integers $X$, we get the distribution function class corresponding to ${\mathcal{C}}_{\phi(x;y_{1},y_{2})}$.
@@
If we reverse the role of parameters and range values, we get a function class indexed by random integers, mapping each random pair of integers to the probability. If we restrict this function class to deterministic pairs, we get the dual distribution function class for $\phi(x;y_{1},y_{2})$.
@@ 5 PAC learning of statistical classes
theoremsupervisedrandomvarclass For a function class ${\mathcal{H}}$ having $\frac{\epsilon}{50}$ fat-shattering dimension at most $d$, one can perform agnostic PAC learning on the randomization class of ${\mathcal{H}}$ with sample complexity:
@@
If ${\mathcal{H}}$ is $\{0,1\}$-valued and has VC-dimension at most $d$, one can perform agnostic PAC learning on the randomization class of ${\mathcal{H}}$ with sample complexity:
@@
Thus we get the same bounds for the distribution class and (moving to the dimension of the dual class) for the dual distribution class.
Example 5.1. Recall Example 1, where we consider a family ${\mathcal{H}}$ of rational functions definable by a partitioned formula $\phi(x,y;\vec{p})$ in the real field, which can be considered as real-valued formula $\tau(x;\vec{p})$ The distribution function class of ${\mathcal{H}}$ is parameterized by random $\vec{p}$. It is agnostic PAC learnable with sample complexity as in Theorem 5. Thus given supervision based on the expectations for various $x_{i}$, we can learn the hypothesis in the distribution class that has the best expected fit in terms of sum of differences.
Note that the base class is defined by equalities over the reals, so it is a restriction of a function class defined by equalities over the complex field. Thus our later results (Section 6) will imply that the statistical classes are also learnable in an online setting.
@@ 5.1 Combinatorial and statistical tools
We will follow the methods of (Ben Yaacov, 2009), which relate upper and lower bound on the combinatorial fat-shattering dimensions to upper and lower bounds on the Rademacher mean width.
Definition 9.
[Height of a set] Let $A\subseteq{\mathcal{R}}^{n}$ be bounded. We define $h_{A}:{\mathcal{R}}^{n}\to{\mathcal{R}}$, the height of $A$ in a particular direction, to be
$h_{A}(\vec{b})=\sup_{\vec{a}\in A}\vec{a}\cdot\vec{b}.$
Lemma 2.
For any bounded $A\subseteq{\mathcal{R}}^{n}$, the height function $h_{A}:{\mathcal{R}}^{n}\to{\mathcal{R}}$ is Borel measurable.
Proof.
If $A$ is countable, then $h_{A}$ is the supremum of a countable family of continuous functions, and is thus Borel measurable.
If $D\subseteq A$, then clearly $h_{D}\leq h_{A}$. If $D$ is dense in $A$, then for every $\vec{a}\in A$, $\vec{v}\in{\mathcal{R}}^{n}$, and $\epsilon>0$, there is some $\vec{d}\in D$ such that $|\vec{v}\cdot(\vec{a}-\vec{d})|\leq\epsilon$, so we can conclude that $h_{D}(\vec{v})\geq h_{A}(\vec{v})-\epsilon$, so in fact, $h_{D}=h_{A}$.
Every subspace of ${\mathcal{R}}^{n}$ is separable, so for every $A$ there is a countable dense $D$, and $h_{A}=h_{D}$ is measurable. ∎
Definition 10.
[Mean width] Let $\beta$ be a Borel probability measure on ${\mathcal{R}}^{n}$. Define the mean width of $A$ w.r.t. $\beta$, $w(A,\beta)$, as $\mathbb{E}_{\beta}\left[h_{A}(\vec{b})\right]$, where $\vec{b}$ is a random variable with distribution $\beta$. By the measurability result mentioned above, this is always defined.
If $\beta$ is the distribution that samples uniformly from $\{+1,-1\}$ $n$ times independently, we define the Rademacher mean width, denoted $w_{\mathcal{R}}(A)$, as $w(A,\beta)$.
Let $f(x,y)$ be a function on $X\times Y$, we let $f(\bar{x},b)$ for $\bar{x}=(x_{1},\dots,x_{n})\in X^{n}$ denote the vector $(f(x_{1};b),\dots,f(x_{n};b))$, and let $f(\bar{x},Y)$ be the set of vectors $\{f(\bar{x},b):b\in Y\}$.
Then we extend the definition of Rademacher mean width to be a function of an integer $n$, given the function $f$; ${\mathcal{R}}_{f}(n)=\sup_{\bar{x}\in X^{n}}w_{\mathcal{R}}(f(\bar{x},Y))$. We will refer to this function as the Rademacher mean width of the function $f$, but this only differs from the “Rademacher complexity” of $f$ in other literature by a factor of $n$.
The reason for looking at Rademacher mean width will be that we can show it behaves well under averaging with respect to an arbitrary measure: see Theorem 1 to follow.
From a bound on the Rademacher width of a function class, we can infer a bound on the Rademacher width of its expectation. Using this and some relationships between Rademacher width bounds and fat-shattering, we are able to bootstrap from a class to its randomization, proving Theorem 5.
Glivenko-Cantelli Dimension. Our arguments for PAC learnability will go through the following dimension:
Definition 11 (Glivenko-Cantelli dimension).
@@
$GC_{{\mathcal{H}}}(\epsilon,\delta)=\min~{}\{n:\forall m\geq n,\forall D\mbox{ Distribution on }X~{}$
$D^{m}\left\{(x_{1},...,x)~{}|~{}\forall h\in{\mathcal{H}},\left|\frac{1}{m}\cdot(\Sigma^{m}_{i=1}h(x_{i}))-\int h(u)dD(u)\right|>\epsilon\right\}\leq\delta\}$
Recall that the law of large numbers implies that if we fix any bounded measurable function $f$ and are given a $\delta$ and $\epsilon$, then we can find an $n$ so that, for any distribution $D$, for all but $\delta$ of the $n$-samples from the distribution, the sample mean of $f$ is within $\epsilon$ of the the mean of $f$. Most proofs that a class is agnostic PAC learnable go through showing that for each $\epsilon,\delta>0$, the dimension $GC_{{\mathcal{H}}}(\epsilon,\delta)$ is finite. From the Glivenko-Cantelli dimension for $\epsilon$ and $\delta$, we can easily obtain bounds on the number of samples needed to learn to given tolerances $\delta$ and $\epsilon$.
Glivenko-Cantelli bounds are used to derive learnability bounds in (Alon et al., 1997) and (Bartlett & Long, 1995, Theorem 14). The proof of (Anthony & Bartlett, 2009, Theorem 19.1) gives the following version of the connection between these concepts:
Fact 11 (Anthony & Bartlett (2009)).
@@
Because the bounds we will derive for $GC_{\mathcal{H}}(\varepsilon,\delta)$ will always be polynomial in $\varepsilon$ and $\delta$, the factors of 2 in this fact will only change the bound by a constant multiple, which will only change the constant of asymptotic notation. While we will need to prove Glivenko-Cantelli bounds for most classes we study, it will be helpful to refer to the following general bound, which, combined with a version of Fact 11, comprised the proof of (Bartlett & Long, 1995, Theorem 14)(our Fact 4):
Fact 12 (Bartlett & Long (1995, Theorem 9)).
The Glivenko-Cantelli dimension is bounded by
$GC_{\mathcal{H}}(\varepsilon,\delta)=O\left({\frac{1}{\epsilon^{2}}}\cdot\left(\mathrm{FatSHDim}_{\frac{\epsilon}{9}}\left({\mathcal{H}}\right)\cdot\log^{2}\left(\frac{1}{\epsilon}\right)+\log\left(\frac{1}{\delta}\right)\right)\right).$

### v2

@@ 3.1 Expectation of a Measurable Family of Hypothesis Classes
We will show that we can embed the distribution class and the dual distribution class in a more general construction for hypothesis classes, where parameters and range elements are treated symmetrically. This construction is fundamentally connected to the randomization construction for hypothesis classes coming from logical formulas, originating with (Keisler, 1999), refined later in (Ben Yaacov and Keisler, 2009; Ben Yaacov, 2009).
We will start with a construction that works on something much more general than a hypothesis class.
Definition 5 (Measurable family of hypothesis classes)
Assume that $(\Omega,\Sigma,\mu)$ is a probability space, and ${\mathcal{F}}=({\mathcal{H}}_{\omega}:\omega\in\Omega)$ is a family of hypothesis classes ${\mathcal{H}}_{\omega}:X\times Y\to[0,1]$ such that for each $x\in X,y\in Y$, the function $\omega\mapsto{\mathcal{H}}_{\omega}(x,y)$ is measurable. We call such a family a measurable family of hypothesis classes on $X$ parameterized by $Y$.
Thus a measurable family can be thought of as a randomized hypothesis class, where for each $x\in X,y\in Y$ we randomly choose a real-valued output.
Definition 6 (Expectation class of a measurable family of hypothesis classes)
Let ${\mathcal{F}}=({\mathcal{H}}_{\omega}:\omega\in\Omega)$ be a measurable family of hypothesis classes. Then the expectation class of ${\mathcal{F}}$ is the function ${\mathbb{E}}{\mathcal{F}}:X\times Y\to[0,1]$ defined by
${\mathbb{E}}{\mathcal{F}}(x,y)=\mathbb{E}_{\mu}[{\mathcal{H}}_{\omega}(x,y)].$
This is a single hypothesis class, with range space $X$ and parameter space $Y$.
3.2 Embedding Distribution Classes in Expectation Classes
Definition 7 ($X$-valued random variables)
@@ 3.2 Embedding Distribution Classes in Expectation Classes
If $(\Omega,\Sigma,\mu)$ is a probability space and $(X,\Sigma_{X})$ is a measurable space, then let ${\mathcal{RV}}_{X}$ be the set of $X$-valued random variables over $\Omega$: that is, measurable functions from $\Omega\to X$.
Definition 8 (Compatibility with a hypothesis class; Randomized version of a hypothesis class)
Let ${\mathcal{H}}$ be a hypothesis class on range space $X$ parameterized by $Y$, and let $\Omega^{*}=(\Omega,\Sigma,\mu)$ be a probability space. A $Y$-valued random variable $Y^{\prime}$ is said to be compatible with ${\mathcal{H}}$ if for each $x\in X$ the function mapping $\omega\in\Omega$ to $h_{Y^{\prime}(\omega)}(x)$ is measurable in the usual sense.
We write $\mathrm{CompatParamRV}({\mathcal{H}})$ for the set of $Y$-valued random variables that are compatible with a given ${\mathcal{H}}$, omitting the dependence on $\Omega^{*}$ for brevity.
For a given $Y^{\prime}\in\mathrm{CompatParamRV}({\mathcal{H}})$ we refer to the real-valued function on $X$ above as the $Y^{\prime}$-randomized version of ${\mathcal{H}}$.
Thus each compatible $Y$-valued random variables thus give an alternative notion of “random parameter”.
Definition 9 (Parameter randomized version of a hypothesis class)
Given a hypothesis class ${\mathcal{H}}$ on range space $X$ parameterized by $Y$ and probability space $(\Omega,\Sigma,\mu)$, we define a measurable family, which we will call the parameter randomized family of ${\mathcal{H}}$, denoted $\mathrm{ParamRandom}({\mathcal{H}})=((\mathrm{ParamRandom}({\mathcal{H}}))_{\omega}:\omega\in\Omega)$. Given $\omega\in\Omega$, let $(\mathrm{ParamRandom}({\mathcal{H}}))_{\omega}$ be the hypothesis class on $X$ consisting of the functions: For each $Y^{\prime}\in\mathrm{CompatParamRV}({\mathcal{H}})$, the function $\lambda x\in X.h_{Y^{\prime}(\omega)}(x)$.
Finally, we can define the expectation class that we want:
Definition 10 (Parameter randomized expectation class)
We consider the hypothesis class denoted ${\mathcal{H}}(\mathrm{CompatParamRV}({\mathcal{H}}))$: it is on the same range space $X$, but parameterized by $\mathrm{CompatParamRV}({\mathcal{H}})$. Given $Y^{\prime}\in\mathrm{CompatParamRV}({\mathcal{H}})$ $h_{Y^{\prime}}$ maps $x\in X$ to $E_{\mu}$ of the $Y^{\prime}$-randomized version of ${\mathcal{H}}$. That is, it is the expectation class of the parameter randomized family of ${\mathcal{H}}$.
The parameter randomized expectation class of a hypothesis class ${\mathcal{H}}$ is very similar to the distribution class of ${\mathcal{H}}$. The difference is that the former considers $Y$-valued functions from a fixed probability space, while the latter deals with the induced distributions on $Y$, ignoring the underlying sample space. We now show that by choosing the probability space carefully, we can close the gap.
Let $\nu$ be a distribution on $Y$, where the underlying $\Sigma$-algebra is generated by $S_{x,I}=\{p\in Y|h_{p}(x)\in I\}$ for $I$ an interval and $x\in X$. We say $\nu$ is induced by a function $f$ from $\Omega$ to $Y$ (with respect to $(\Omega,\Sigma,\mu)$) if $\nu(S_{x,I})=\mu(\{\omega|f(\omega)\in S_{x,I}\}$.
The following simple result in measure theory states that by choosing $(\Omega,\Sigma,\mu)$ appropriately, all the distributions over $Y$ are induced by functions from $(\Omega,\Sigma,\mu)$:
Proposition 11
For any $Y$ we can find $(\Omega,\Sigma,\mu)$ such that for every distribution $\nu$ over the $\Sigma$-algebra on $Y$ generated by $S_{x,I}$, there is a function $f$ such that $\nu$ is induced by $f$ with respect to $(\Omega,\Sigma,\nu)$.
Although the result is well-known (see, e.g. Fremlin (2002)), we sketch the proof:
Proof  We choose $\Omega^{*}=(\Omega,\Sigma,\mu)$ to be a product of all measure spaces on $Y$. Then every measure is induced as a projection on one component.
From the proposition we see:
Proposition 12
By choosing $\Omega^{*}$ as in Proposition 11, we have ${\mathcal{H}}(\mathrm{CompatParamRV}({\mathcal{H}}))$ contains all the functions in the distribution class of ${\mathcal{H}}$. Thus if the former is learnable, in any of the variants we have discussed (agnostic PAC, agnostic online etc.), then so is the latter, with the same sample complexity bounds.
We can analogously talk about $X$-valued random variable $X^{\prime}$ being compatible with ${\mathcal{H}}$, and define the class ${\mathcal{H}}(\mathrm{CompatRangeRV}({\mathcal{H}}))$ of hypothesis parameterized by such random variables: these subsume the dual distribution class.
We will use these embeddings below to transfer results about learnability of the expectation class of a measurable family of hypothesis classes to results about the distribution and dual distribution classes. We emphasize that the notion of measurable family is much more general than the distribution class and dual distribution class. In the distribution class, only parameters are randomized, while range elements are deterministic, while in the dual distribution class it is the reverse. The measurable family allows us to model situations where the range elements and parameters are both randomized, in a correlated way.
Example 3.3. Consider a concept class ${\mathcal{C}}$ containing characteristic functions $h(x)_{y_{1},y_{2}}$ where both the range elements $x$ and the parameters $y_{1},y_{2}$ are integers. The concept contains any $x$ that an even integer between $y_{1}$ and $y_{2}$. Thus as we vary the parameters $y_{1},y_{2}$ we get the set of intervals intersected with the even numbers.
Fix a probability space $\Omega^{*}=(\Omega,\Sigma,\mu_{0})$. The parameterized randomized family of ${\mathcal{H}}$ is a measuable family of hypothesis class parameterized by a “random pair of integers”: a measurable function taking $\omega\in\Omega$ to a pair $y_{1},y_{2}$. The measurable family contains a class for each $\omega\in\Omega$ parameterized by such functions $P$, where for each $\omega$ and $P$, the function maps $x$ to $C(x)_{P(\omega)}$. The expectation class of this measurable family is the distribution class of the concept class ${\mathcal{C}}$.
@@
If we reverse the role of parameters and range values, we get a measurable family of concept classes indexed by random integers $X$, where given $\omega$ and $X$ we map $(y_{1},y_{2})$ to $C(X(\omega))_{y_{1},y_{2}}$. The expectation class of this measurable family is the dual distribution class of ${\mathcal{C}}$
We could also consider measurable families parameterized by functions from $\omega$ to triples $(x,y_{1},y_{2})$, which represent a correlated pair $X,P$ of random range elements and random parameters.
While the notation used in Ben Yaacov (2009) is different, it is proven there that a uniform bound on fat-shattering dimensions of the hypothesis classes in a measurable family implies finite fat-shattering dimension, and thus PAC learnability, of the expectation class.
Definition 13
For $\gamma>0$, say that a measurable family of hypothesis classes ${\mathcal{F}}$ has uniformly bounded $\gamma$ fat-shattering dimension it there is some $d$ such that each class ${\mathcal{H}}\in{\mathcal{F}}$ has $\gamma$ fat-shattering dimension at most $d$.
Fact 9 (Ben Yaacov (2009, Corollaries 4.2, 4.3))
If ${\mathcal{F}}$ is a measurable family of hypothesis classes, and for every $\gamma>0$, the family has uniformly bounded $\gamma$ fat-shattering dimension, then ${\mathbb{E}}{\mathcal{F}}$ has finite fat-shattering dimension.
Now let us return to the distribution class, and recall that it can be embedded in the parameter randomized family of ${\mathcal{H}}$, which associates the function $\lambda x\in X.h_{Y^{\prime}(\omega)}(x)$. If ${\mathcal{H}}$ has finite $\gamma$ fat-shattering dimension $d$, then for any $\omega$, the corresponding class of functions above also has fat-shattering dimension $d$: thus the measurable family has uniformly bounded $\gamma$ fat-shattering dimension. Similarly for the dual distribution class. Thus we get the following corollary of Fact 9:
Corollary 14
If ${\mathcal{H}}$ is agnostic PAC learnable, so are the distribution class and dual distribution classes.
In Section 4, we will adapt the proof from Ben Yaacov (2009) to provide concrete sample complexity bounds on learnability of the expectation class of a measurable family, which will provide improved bounds for the distribution and dual distribution class.
@@ 4 PAC learning of statistical classes
Remark 15
A warning that throughout this section and the next, we ignore several measurability issues that arise when uncountable hypothesis classes are considered – e.g. measurability of sets that involve intersecting over uncountabiy many objects. These subtleties do not arise when the parameter set and range sets are countable, and all the results in the next sections are true without qualification in this case. Extensions to the uncountable case require further sanity conditions on the hypothesis classes. We defer a discussion of this to Appendix C.
@@
theoremsupervisedrandomvarclass If ${\mathcal{F}}$ is a measurable family of hypothesis classes such that each class ${\mathcal{H}}\in{\mathcal{F}}$ has $\frac{\epsilon}{50}$ fat-shattering dimension at most $d$, one can perform agnostic PAC learning on the expectation class ${\mathbb{E}}{\mathcal{F}}$ with sample complexity:
@@
If ${\mathcal{F}}$ is a measurable family of $\{0,1\}$-valued concept classes with VC-dimension at most $d$, one can perform agnostic PAC learning on the expectation class ${\mathbb{E}}{\mathcal{F}}$ with sample complexity:
@@
Thus, via Proposition 12, we get the same bounds for the distribution class and (moving to the dimension of the dual class) for the dual distribution class.
Example 4.1. Recall Example 4, where we consider a family ${\mathcal{H}}$ of rational functions definable by real-coefficients $\vec{p}$. The distribution function class of ${\mathcal{H}}$ is parameterized by random $\vec{p}$. It is agnostic PAC learnable with sample complexity as in Theorem 15. Thus given supervision based on the expectations for various $x_{i}$, we can learn the hypothesis in the distribution class that has the best expected fit in terms of sum of differences.
Thus the next few subsections will be devoted to the proof of Theorem 15. which will complete our sample complexity analysis of agnostic PAC learning for statistical classes.
@@ 4.1 Combinatorial and statistical tools
Our arguments for PAC learnability will go through the following dimension:
Definition 16 (Glivenko-Cantelli dimension)
@@
$GC_{{\mathcal{H}}}(\epsilon,\delta)=\min\leavevmode\nobreak\ \{n:\forall m\geq n,\forall D\mbox{ Distribution on }X\leavevmode\nobreak$
$D^{m}\left\{(x_{1},...,x_{m})\leavevmode\nobreak\ |\leavevmode\nobreak\ \exists h\in{\mathcal{H}},\left|\frac{1}{m}\cdot(\Sigma^{m}_{i=1}h(x_{i}))-\int h(u)dD(u)\right|>\epsilon\right\}\leq\delta\}$
Recall that the law of large numbers implies that if we fix any bounded measurable function $f$ into the reals, and are given a $\delta$ and $\epsilon$, then we can find an $n$ so that, for any distribution $D$, for all but $\delta$ of the $n$-samples from the distribution, the sample mean of $f$ is within $\epsilon$ of the the mean of $f$. Most proofs that a class is agnostic PAC learnable go through showing that for each $\epsilon,\delta>0$, the dimension $GC_{{\mathcal{H}}}(\epsilon,\delta)$ is finite. From the Glivenko-Cantelli dimension for $\epsilon$ and $\delta$, we can easily obtain bounds on the number of samples needed to learn to given tolerances $\delta$ and $\epsilon$.
Glivenko-Cantelli bounds are used to derive learnability bounds in (Alon et al., 1997) and (Bartlett and Long, 1995, Theorem 14). The proof of (Anthony and Bartlett, 2009, Theorem 19.1) gives the following version of the connection between these concepts:
Fact 10 (Anthony and Bartlett (2009))
@@
Because the bounds we will derive for $GC_{\mathcal{H}}(\varepsilon,\delta)$ will always be polynomial in $\varepsilon$ and $\delta$, the factors of $2$ in this fact will only change the bound by a constant multiple, which will only change the constant within asymptotic notation.
We will follow the methods of (Ben Yaacov, 2009), which relate upper and lower bound on the combinatorial fat-shattering dimensions to upper and lower bounds on the Rademacher mean width. We will be able to then move to bounds on the Rademacher width of the expectation class, and from there to bound on GC-dimension of the expection class.
Definition 17
[Width of a set in a direction] Let $A\subseteq{\mathcal{R}}^{n}$ be bounded. Given $\vec{b}\in{\mathcal{R}}^{n}$, we define $0pt(A,\vec{b})$, the width of $A$ in the particular direction $\vec{b}$, to be $0pt(A,\vec{b})=\sup_{\vec{a}\in A}\vec{a}\cdot\vec{b}.$
Lemma 18
For any bounded $A\subseteq{\mathcal{R}}^{n}$, the width function $\vec{b}\mapsto 0pt(A,\vec{b})$ is Borel measurable.
Proof  If $A$ is countable, then $0pt(A,\vec{b})$ is the supremum of a countable family of continuous functions of $\vec{b}$, and is thus Borel measurable.
If $D\subseteq A$, then for every $\vec{b}$, $0pt(D,\vec{b})\leq 0pt(A,\vec{b})$. If $D$ is dense in $A$, then for every $\vec{a}\in A$, $\vec{b}\in{\mathcal{R}}^{n}$, and $\epsilon>0$, there is some $\vec{d}\in D$ such that $|\vec{b}\cdot(\vec{a}-\vec{d})|\leq\epsilon$, so we can conclude that $0pt(D,\vec{b})\geq 0pt(A,\vec{b})-\epsilon$, so in fact, $0pt(D,\vec{b})=0pt(A,\vec{b})$.
Every subspace of ${\mathcal{R}}^{n}$ is separable, so for every $A$ there is a countable dense $D$, and $\vec{b}\mapsto 0pt(D,\vec{b})$, and thus $\vec{b}\mapsto 0pt(A,\vec{b})$, is measurable.
Definition 19
[Mean width] Let $\beta$ be a Borel probability measure on ${\mathcal{R}}^{n}$. Define the mean width of $A$ w.r.t. $\beta$, $w(A,\beta)$, as $\mathbb{E}_{\beta}\left[0pt(A,\vec{b})\right]$, where $\vec{b}$ is a random variable with distribution $\beta$. By the measurability result mentioned above, this is always defined.
If $\beta$ is the distribution that samples uniformly from $\{+1,-1\}^{n}$, we define the Rademacher mean width, denoted $w_{\mathcal{R}}(A)$, as $w(A,\beta)$.
For any function $h:X\to[0,1]$ and $\bar{x}=(x_{1},\dots,x_{n})\in X^{n}$, we let $h(\bar{x})$ denote the vector $(h(x_{1}),\dots,h(x_{n}))$. In particular, if ${\mathcal{H}}$ is a hypothesis class on $X$ parameterized by $Y$, $\bar{x}=(x_{1},\dots,x_{n})\in X^{n}$, and $y\in Y$, then ${\mathcal{H}}(\bar{x},y)=({\mathcal{H}}(x_{1},y),\dots,{\mathcal{H}}(x_{n},y))$.
For any hypothesis class ${\mathcal{H}}$ on $X$ parametrized by $Y$ and $\bar{x}=(x_{1},\dots,x_{n})\in X^{n}$, let ${\mathcal{H}}(\bar{x},Y)$ be the set of vectors
$\{h(\bar{x}):h\in{\mathcal{H}}\}=\{{\mathcal{H}}(\bar{x},y):y\in Y\}\subseteq[0,1]^{n}.$
Then we extend the definition of Rademacher mean width to be a function of an integer $n$, given the class ${\mathcal{H}}$ on $X$ parametrized by $Y$; ${\mathcal{R}}_{\mathcal{H}}(n)=\sup_{\bar{x}\in X^{n}}w_{\mathcal{R}}({\mathcal{H}}(\bar{x},Y))$. We will refer to this function as the Rademacher mean width of the class ${\mathcal{H}}$, but this only differs from the “Rademacher complexity” of ${\mathcal{H}}$ in other literature by a factor of $n$.
The reason for looking at Rademacher mean width will be that it behaves well under averaging with respect to an arbitrary measure: see Theorem 22 to follow.
From a bound on the Rademacher width of a function class, we can infer a bound on the Rademacher width of its expectation. Using this and some relationships between Rademacher width bounds and fat-shattering, we are able to bootstrap from uniform bounds on a measurable family to bounds on the expectation class, proving Theorem 15.